# SDEs with supercritical distributional drifts

### v1

@@ 3.1. Subcritical case: $\frac{d}{p}+\frac{2}{q}<1+\alpha$
    Let $(\alpha_{b},p_{b},q_{b})\in(-1,-\frac{1}{2}]\times[2,\infty]^{2}$ with $\frac{d}{p_{b}}+\frac{2}{q_{b}}<1+\alpha_{b}$. Suppose that
    $\kappa^{b}_{1}:=\|b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}<\infty\quad\text{and}\quad\kappa^{b}_{2}:=\|\mathord{{\rm div}}b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{-2-\alpha_{b}}_{p_{b},q_{b}/(q_{b}-1)}}<\infty.$
@@
$2+\alpha_{b}+\tfrac{d}{p_{b}}-\tfrac{2}{q_{b}}<2-\tfrac{4}{q_{b}}\leqslant 2-\tfrac{2}{q}.$ (3.2)
@@
$\kappa^{b}_{1}=\|b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{-1/2}_{p_{b},q_{b}}}\lesssim\|b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{-1/2}_{p_{b},1}},\ \ \kappa^{b}_{2}=\|\mathord{{\rm div}}b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{-3/2}_{p_{b},q_{b}/(q_{b}-1)}}\lesssim\|b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{-1/2}_{p_{b},1}}.$
If $b$ is time-independent, then it was covered by the condition $b\in{\mathbf{H}}^{-1}_{p_{b}}$ with $p_{b}>2d$ in [48].
@@
$\Theta:=(T,d,\alpha_{b},p_{b},q_{b},\kappa^{b}_{1},\kappa^{b}_{2}).$ (3.3)
@@
For any $f\in{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}$, there is a $\lambda_{0}=\lambda_{0}(\Theta)>0$ such that for any $\lambda\geqslant\lambda_{0}$, there is a unique $u^{\lambda}$ with the regularity that for any $\theta\in(0,2+\alpha_{b}+\frac{d}{p_{b}}-\frac{2}{q_{b}}]$,
$\|u^{\lambda}\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{2+\alpha_{b}}_{p_{b},q_{b}}}+(1+\lambda)^{\frac{\theta}{2}}\|u^{\lambda}\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{2+\alpha_{b}-2/q_{b}-\theta}_{p_{b},1}}\leqslant C\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}},$ (3.4)
where $C=C(\Theta,\alpha,p,q)>0$, which solves the following integral equation:
$u^{\lambda}_{t}={\mathscr{I}}^{\lambda}_{t}(b\odot\nabla u^{\lambda}-\mathord{{\rm div}}b\preceq u^{\lambda}+f),$
where ${\mathscr{I}}^{\lambda}$ is defined by (2.18).
@@
We only show the a priori estimate (3.4). The existence follows by standard Picard’s iteration. For simplicity of notations, we drop the superscript $\lambda$ over $u^{\lambda}$.
By Schauder’s estimate (2.22) and (2.16), we have for $\theta\in(0,2-\frac{2}{q_{b}})$,
$\begin{split}&\|{\mathscr{I}}^{\lambda}(b\odot\nabla u)\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{2+\alpha_{b}}_{p_{b},q_{b}}}+(1+\lambda)^{\frac{\theta}{2}}\|{\mathscr{I}}^{\lambda}(b\odot\nabla u)\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{2+\alpha_{b}-2/q_{b}-\theta}_{p_{b},1}}\\ &\qquad\lesssim_{C}\|b\odot\nabla u\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}\lesssim_{C}\kappa^{b}_{1}\|u\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{1+d/p_{b}}_{p_{b},1}},\end{split}$ (3.5)
where $C=C(\Theta,\theta)>0$. On the other hand, by (2.21), we have for any $\theta^{\prime}\in(0,2-\frac{2}{q_{b}})$,
$\|{\mathscr{I}}^{\lambda}(\mathord{{\rm div}}b\preceq u)\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{2-d/p_{b}-2/q_{b}-\theta^{\prime}}_{p_{b},1}}\lesssim_{C}(1+\lambda)^{-\frac{\theta^{\prime}}{2}}\|\mathord{{\rm div}}b\preceq u\|_{{\mathbb{L}}^{q_{b}/2}_{T}{\mathbf{B}}^{0}_{p_{b}/2,\infty}},$ (3.6)
@@
$\|{\mathscr{I}}^{\lambda}(\mathord{{\rm div}}b\preceq u)\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{2-d/p_{b}-2/q_{b}-\theta^{\prime\prime}}_{p_{b},1}}\lesssim_{C}(1+\lambda)^{-\frac{\theta^{\prime\prime}}{2}}\|\mathord{{\rm div}}b\preceq u\|_{{\mathbb{L}}^{q_{b}/2}_{T}{\mathbf{B}}^{0}_{p_{b}/2,\infty}}.$ (3.7)
@@
If we take $\theta^{\prime}=\vartheta$ in (3.6) and $\theta^{\prime\prime}=\vartheta+\theta$ in (3.7), then by (2.17),
$\begin{split}&\|{\mathscr{I}}^{\lambda}(\mathord{{\rm div}}b\preceq u)\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{2+\alpha_{b}}_{p_{b},1}}+(1+\lambda)^{\frac{\theta}{2}}\|{\mathscr{I}}^{\lambda}(\mathord{{\rm div}}b\preceq u)\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{2+\alpha_{b}-2/q_{b}-\theta}_{p_{b},1}}\\ &\qquad\lesssim_{C}(1+\lambda)^{-\frac{\vartheta}{2}}\|\mathord{{\rm div}}b\preceq u\|_{{\mathbb{L}}^{q_{b}/2}_{T}{\mathbf{B}}^{0}_{p_{b}/2,\infty}}\lesssim_{C}(1+\lambda)^{-\frac{\vartheta}{2}}\kappa^{b}_{2}\|u\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{2+\alpha_{b}}_{p_{b},q_{b}}},\end{split}$ (3.8)
where $C=C(\Theta,\theta,\vartheta)>0$. Moreover, by $\alpha-\tfrac{d}{p}-\tfrac{2}{q}\geqslant\alpha_{b}-\tfrac{d}{p_{b}}-\tfrac{2}{q_{b}}$ and (2.22), for $\theta\in(0,2-\frac{2}{q})$,
$\|{\mathscr{I}}^{\lambda}(f)\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{2+\alpha_{b}}_{p_{b},q_{b}}}+(1+\lambda)^{\frac{\theta}{2}}\|{\mathscr{I}}^{\lambda}(f)\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{2+\alpha_{b}-2/q_{b}-\theta}_{p_{b},1}}\lesssim_{C}\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}}.$ (3.9)
Combining (3.5), (3.8), (3.9) and by (3.2), we obtain that for any $\theta\in(0,2+\alpha_{b}+\frac{d}{p_{b}}-\tfrac{2}{q_{b}}]$ and $\vartheta\in(0,\vartheta_{0})$,
$\begin{split}&\|u\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{2+\alpha_{b}}_{p_{b},q_{b}}}+(1+\lambda)^{\frac{\theta}{2}}\|u\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{2+\alpha_{b}-2/q_{b}-\theta}_{p_{b},1}}\\ &\qquad\leqslant C_{1}\kappa^{b}_{1}\|u\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{1+d/p_{b}}_{p_{b},1}}+C_{2}(1+\lambda)^{-\frac{\vartheta}{2}}\kappa^{b}_{2}\|u\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{2+\alpha_{b}}_{p_{b},q_{b}}}+C_{3}\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}},\end{split}$ (3.10)
where $C_{1}=C_{1}(\Theta,\theta)>0$, $C_{2}=C_{2}(\Theta,\theta,\vartheta)>0$ and $C_{3}=C_{3}(\Theta,\theta,\alpha,p,q)>0$.
Since $\frac{2}{q_{b}}+\frac{d}{p_{b}}<1+\alpha_{b}$, one can choose $\theta_{0}=\theta_{0}(d,\alpha_{b},p_{b},q_{b})$ being close to zero so that
$\|u\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{1+d/p_{b}}_{p_{b},1}}\lesssim_{C}\|u\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{2+\alpha_{b}-2/q_{b}-\theta_{0}}_{p_{b},1}}.$
@@
$\|u\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{2+\alpha_{b}}_{p_{b},q_{b}}}+(1+\lambda)^{\frac{\theta_{0}}{2}}\|u\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{2+\alpha_{b}-2/q_{b}-\theta_{0}}_{p_{b},1}}\lesssim_{C}\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}}.$
@@ 3.2. Supercritical case: $\frac{d}{p}+\frac{2}{q}<2+\alpha$
${\mathscr{I}}_{d}:=\Big{\{}(\alpha,p,q)\in[-1,0]\times[2,\infty]^{2},\ \tfrac{d}{p}+\tfrac{2}{q}<2+\alpha\Big{\}},$ (3.11)
@@
${\mathscr{V}}_{T}:=\left\{u:\|u\|_{{\mathscr{V}}_{T}}:=\sup_{t\in[0,T]}\|u(t)\|_{2}+\|\nabla u\|_{{\mathbb{L}}^{2}_{T}}<\infty\right\}.$
@@
For any $(\alpha,p,q)\in{\mathscr{I}}_{d}$ and $T>0$, there is a constant $C=C(\alpha,p,q,d,T)>0$ such that for any $f\in{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}$ and $u,g\in{\mathscr{V}}_{T}$,
$\|\langle fu,g\rangle\|_{L^{1}_{T}}\lesssim_{C}\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}}\|u\|_{{\mathscr{V}}_{T}}\|g\|_{{\mathscr{V}}_{T}}.$
@@
Let $(\alpha,p,q)\in{\mathscr{I}}_{d}$. Consider the case $p\in[2,\infty)$. Let $r\in(1,2]$ be defined by $\frac{1}{p}+\frac{1}{r}=1$. Let us choose $r_{0}\in(1,r]$ so that $-\alpha-\frac{d}{r}\leqslant 1-\frac{d}{r_{0}}$. By Sobolev’s embedding (2.9), we have
@@
Let $\frac{1}{r_{0}}=\frac{1}{r_{1}}+\frac{1}{2}$. By (2.5) and Hölder’s inequality, we have
@@
Since $p,q\in[2,\infty)$ satisfies $\frac{d}{p}+\frac{2}{q}<2+\alpha$, one has $r_{1},q_{1}\in[2,\infty)$ and $\frac{d}{r_{1}}+\frac{2}{q_{1}}>\frac{d}{2}$. By [49, Lemma 2.1], we conclude the desired estimate.
@@
    Let $(\alpha_{b},p_{b},q_{b})\in{\mathscr{I}}_{d}$. Suppose that
    $\kappa^{b}_{1}:=\|b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}}<\infty\quad\text{and}\quad\kappa^{b}_{2}:=\|\mathord{{\rm div}}b\|_{{\mathbb{L}}^{2}_{T}L^{\infty}}<\infty.$
Recall the parameter set (3.3). Now we can show the following existence and uniqueness result.
@@
$\|u\|_{{\mathbb{L}}^{\infty}_{T}}+\|u\|_{{\mathscr{V}}_{T}}\lesssim_{C}\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}}.$ (3.13)
@@
$\kappa^{b_{n}}_{1}\leqslant\kappa^{b}_{1},\ \ \kappa^{b_{n}}_{2}\leqslant\kappa^{b}_{2}.$

### v3

@@ 3.1. Subcritical case: $\frac{d}{p}+\frac{2}{q}<1+\alpha$
    Let $(\alpha_{b},p_{b},q_{b})\in(-1,-\frac{1}{2}]\times[2,\infty]^{2}$ with $\tfrac{d}{p_{b}}+\tfrac{2}{q_{b}}<1+\alpha_{b}$. Suppose that
    $\kappa_{b}:=\|b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}<\infty\quad\text{and}\quad\tilde{\kappa}_{b}:=\|\mathord{{\rm div}}b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{-2-\alpha_{b}}_{p_{b},q_{b}/(q_{b}-1)}}<\infty.$
@@
$2+\alpha_{b}+\tfrac{d}{p_{b}}-\tfrac{2}{q_{b}}<2-\tfrac{4}{q_{b}}\leqslant 2-\tfrac{2}{q}.$
@@
$\kappa_{b}=\|b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{-1/2}_{p_{b},q_{b}}}\lesssim\|b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{-1/2}_{p_{b},1}},\ \ \tilde{\kappa}_{b}=\|\mathord{{\rm div}}b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{-3/2}_{p_{b},q_{b}/(q_{b}-1)}}\lesssim\|b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{-1/2}_{p_{b},1}}.$
If $b$ is time-independent, then it was covered by the condition $b\in{\mathbf{H}}^{-1/2}_{p_{b}}$ with $p_{b}>2d$ in [50].
@@
$\Theta:=(T,d,\alpha_{b},p_{b},q_{b},\kappa_{b},\tilde{\kappa}_{b}).$ (3.2)
@@
$\|u^{\lambda}\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{2+\alpha_{b}}_{p_{b},q_{b}}}+(1+\lambda)^{\frac{\theta}{2}}\|u^{\lambda}\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{2+\alpha_{b}-2/q_{b}-\theta}_{p_{b},1}}\leqslant C(\Theta,\theta,\alpha,p,q)\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}}.$ (3.3)
@@
We only show the a priori estimate (3.3). The existence follows by standard Picard’s iteration. For simplicity of notations, we drop the superscript $\lambda$ over $u^{\lambda}$, and write for $\lambda,\theta\geqslant 0$,
$\|g\|_{{\mathcal{S}}^{\lambda}_{\theta}}:=\|g\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{2+\alpha_{b}}_{p_{b},q_{b}}}+(1+\lambda)^{\frac{\theta}{2}}\|g\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{2+\alpha_{b}-2/q_{b}-\theta}_{p_{b},1}}.$
By Schauder’s estimate (2.22), (2.6) and (2.16), we have for $\theta\in(0,2-\frac{2}{q_{b}})$,
$\|{\mathscr{I}}^{\lambda}(b\odot\nabla u)\|_{{\mathcal{S}}^{\lambda}_{\theta}}\lesssim\|b\odot\nabla u\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}\lesssim\kappa_{b}\|u\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{1}_{\infty,1}}.$ (3.4)
On the other hand, by (2.21), we have for any $\theta^{\prime}\in(0,2-\frac{2}{q_{b}})$,
$\|{\mathscr{I}}^{\lambda}(\mathord{{\rm div}}b\preceq u)\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{2-d/p_{b}-2/q_{b}-\theta^{\prime}}_{p_{b},1}}\lesssim(1+\lambda)^{-\frac{\theta^{\prime}}{2}}\|\mathord{{\rm div}}b\preceq u\|_{{\mathbb{L}}^{q_{b}/2}_{T}{\mathbf{B}}^{0}_{p_{b}/2,\infty}},$ (3.5)
@@
$\|{\mathscr{I}}^{\lambda}(\mathord{{\rm div}}b\preceq u)\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{2-d/p_{b}-4/q_{b}-\theta^{\prime\prime}}_{p_{b},1}}\lesssim(1+\lambda)^{-\frac{\theta^{\prime\prime}}{2}}\|\mathord{{\rm div}}b\preceq u\|_{{\mathbb{L}}^{q_{b}/2}_{T}{\mathbf{B}}^{0}_{p_{b}/2,\infty}}.$ (3.6)
@@
If we take $\theta^{\prime}=\vartheta$ in (3.5) and $\theta^{\prime\prime}=\vartheta+\theta$, and $\theta^{\prime\prime}=\vartheta$ in (3.6), then by $\alpha_{b}<-d/q_{b}-\vartheta$ and (2.17),
$\|{\mathscr{I}}^{\lambda}(\mathord{{\rm div}}b\preceq u)\|_{{\mathcal{S}}^{\lambda}_{\theta}}\lesssim(1+\lambda)^{-\frac{\vartheta}{2}}\|\mathord{{\rm div}}b\preceq u\|_{{\mathbb{L}}^{q_{b}/2}_{T}{\mathbf{B}}^{0}_{p_{b}/2,\infty}}\lesssim(1+\lambda)^{-\frac{\vartheta}{2}}\tilde{\kappa}_{b}\|u\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{2+\alpha_{b}}_{p_{b},q_{b}}},$ (3.7)
Moreover, by $\alpha-\tfrac{d}{p}-\tfrac{2}{q}\geqslant\alpha_{b}-\tfrac{d}{p_{b}}-\tfrac{2}{q_{b}}$, (2.22) and (2.6), for $\theta\in(0,2-\frac{2}{q})$,
$\|{\mathscr{I}}^{\lambda}(f)\|_{{\mathcal{S}}^{\lambda}_{\theta}}\lesssim\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}}.$ (3.8)
Combining (3.4), (3.7), (3.8) and by Remark 3.1, for $\theta\in(0,2+\alpha_{b}+\frac{d}{p_{b}}-\tfrac{2}{q_{b}}]$ and $\vartheta\in(0,\vartheta_{0})$, there are constants $C_{1}=C_{1}(\Theta,\theta)>0$, $C_{2}=C_{2}(\Theta,\theta,\vartheta)>0$ and $C_{3}=C_{3}(\Theta,\theta,\alpha,p,q)>0$ such that
$\|u\|_{{\mathcal{S}}^{\lambda}_{\theta}}\leqslant C_{1}\kappa_{b}\|u\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{1}_{\infty,1}}+C_{2}(1+\lambda)^{-\frac{\vartheta}{2}}\tilde{\kappa}_{b}\|u\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{2+\alpha_{b}}_{p_{b},q_{b}}}+C_{3}\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}}.$ (3.9)
Since $\frac{2}{q_{b}}+\frac{d}{p_{b}}<1+\alpha_{b}$, by (2.8), one can choose $\theta_{0}=\theta_{0}(d,\alpha_{b},p_{b},q_{b})$ being close to zero so that
$\|u\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{1}_{\infty,1}}\lesssim_{C}\|u\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{2+\alpha_{b}-2/q_{b}-\theta_{0}}_{p_{b},1}}.$
@@
$\|u\|_{{\mathcal{S}}^{\lambda}_{\theta_{0}}}=\|u\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{2+\alpha_{b}}_{p_{b},q_{b}}}+(1+\lambda)^{\frac{\theta_{0}}{2}}\|u\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{2+\alpha_{b}-2/q_{b}-\theta_{0}}_{p_{b},1}}\lesssim_{C}\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}}.$
@@ 3.2. Supercritical case: $\frac{d}{p}+\frac{2}{q}<2+\alpha$
${\mathbf{I}}_{d}:=\Big{\{}(\alpha,p,q)\in[-1,0]\times[2,\infty]^{2},\ \tfrac{d}{p}+\tfrac{2}{q}<2+\alpha\Big{\}},$ (3.10)
@@
${\mathscr{V}}_{T}:=\left\{u:\|u\|_{{\mathscr{V}}_{T}}:=\sup_{t\in[0,T]}\|u(t)\|_{2}+\|\nabla u\|_{{\mathbb{L}}^{2}_{T}}<\infty\right\}.$ (3.11)
@@
For any $(\alpha,p,q)\in{\mathbf{I}}_{d}$ and $T>0$, there is a constant $C=C(\alpha,p,q,d,T)>0$ such that for any $f\in{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}$ and $u,g\in{\mathscr{V}}_{T}$,
$\|\langle fu,g\rangle\|_{L^{1}_{T}}\leqslant C\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}}\|u\|_{{\mathscr{V}}_{T}}\|g\|_{{\mathscr{V}}_{T}}.$
@@
Let $(\alpha,p,q)\in{\mathbf{I}}_{d}$. Consider the case $p\in[2,\infty)$. Let $r\in(1,2]$ be defined by $\frac{1}{p}+\frac{1}{r}=1$. Let us choose $r_{0}\in(1,r]$ so that $-\alpha-\frac{d}{r}\leqslant 1-\frac{d}{r_{0}}$. By Sobolev’s embedding (2.9), we have
@@
Let $\frac{1}{r_{0}}=\frac{1}{r_{1}}+\frac{1}{2}$. By Hölder’s inequality, we have
@@
Since $p,q\in[2,\infty)$ satisfies $\frac{d}{p}+\frac{2}{q}<2+\alpha$, one has $r_{1},q_{1}\in[2,\infty)$ and $\frac{d}{r_{1}}+\frac{2}{q_{1}}>\frac{d}{2}$. By [51, Lemma 2.1], we conclude the desired estimate.
@@
    Let $(\alpha_{b},p_{b},q_{b})\in{\mathbf{I}}_{d}$. Suppose that
    $\kappa_{b}:=\|b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}}<\infty\quad\text{and}\quad\tilde{\kappa}_{b}:=\|\mathord{{\rm div}}b\|_{{\mathbb{L}}^{2}_{T}L^{\infty}}<\infty.$
Recall the parameter set (3.2). We can show the following existence and uniqueness result.
@@
$\|u\|_{{\mathbb{L}}^{\infty}_{T}}+\|u\|_{{\mathscr{V}}_{T}}\leqslant C\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}}.$ (3.13)
@@
$\kappa_{b_{n}}\leqslant\kappa_{b},\ \ \tilde{\kappa}_{b_{n}}\leqslant\tilde{\kappa}_{b}.$